OpenStudy (anonymous):

Find the coordinates of corner point A.−x + 2y = 2, x + y = 2, and y = 0.

OpenStudy (eyust707):

okay so we have three lines

OpenStudy (eyust707):

neither of them or parallel so they intersect to form a triangle

OpenStudy (eyust707):

we are looing for the corners of this triangle

OpenStudy (eyust707):

well the corner occurs when the two lines meet

OpenStudy (eyust707):

when the two lines are equal to each other

OpenStudy (eyust707):

two of the corners are easy to find because we can just plug y=0 into the equaitons

OpenStudy (eyust707):

then useing that x that we get we can solce for the y

OpenStudy (eyust707):

would you like me to show you an example

OpenStudy (anonymous):

That would be great. THis has stomped me for hours.

OpenStudy (eyust707):

okay lets look at these two lines: x + y = 2 y = 0

OpenStudy (eyust707):

we want to find the point where they intersect

OpenStudy (eyust707):

well we know for that bottom line y always equals 0

OpenStudy (eyust707):

since y equals zero we need to find the x to satisfy the top equation when y is 0

OpenStudy (eyust707):

so just plug y=0 into the top equation

OpenStudy (eyust707):

what does x then equal?

OpenStudy (anonymous):

0

OpenStudy (eyust707):

nope. x + 0 = 2

OpenStudy (eyust707):

what is x?

OpenStudy (anonymous):

2

OpenStudy (eyust707):

yes so whe y=0 and x=2 those lines meet

OpenStudy (eyust707):

that is one of your corners

OpenStudy (eyust707):

try and do the same thing for these two lines: −x + 2y = 2 y = 0.

OpenStudy (eyust707):

clearly y equals 0

OpenStudy (eyust707):

but what does x equal?

OpenStudy (anonymous):

-2

OpenStudy (eyust707):

correct!

OpenStudy (eyust707):

so the second corner is y=0 x=-2

OpenStudy (eyust707):

now for the last one: −x + 2y = 2 x + y = 2

OpenStudy (eyust707):

youve probably been doing these in class, do you knwo how to solve these?

OpenStudy (anonymous):

These are my downfall. I am not fond of finding these corner points and the problems she always gave use were so difficult.

OpenStudy (eyust707):

well this one is actually pretty easy , once you get the hang of it

OpenStudy (anonymous):

the first time i tried this last part i got a fraction.

OpenStudy (eyust707):

thats not necessarily a bad thing...

OpenStudy (eyust707):

what did you get?

OpenStudy (anonymous):

y=4/3

OpenStudy (eyust707):

correct!

OpenStudy (eyust707):

and whats x?

OpenStudy (anonymous):

really?

OpenStudy (eyust707):

yep

OpenStudy (anonymous):

omg i stopped there because i just knew i was wrong

OpenStudy (eyust707):

nope not at all!

OpenStudy (anonymous):

i can put the y into either equation, right?

OpenStudy (eyust707):

yep

OpenStudy (anonymous):

x=2/3

OpenStudy (eyust707):

thats it so your third point is : y=4/3 x=2/3

OpenStudy (eyust707):

(2/3 , 4/3)

OpenStudy (anonymous):

thank you so much! :)